# TrojanRAG: Retrieval-Augmented Generation Can Be Backdoor Driver in Large Language Models

### v2

@@ Abstract
Large language models (LLMs) have raised concerns about potential security threats despite performing significantly in Natural Language Processing (NLP). Backdoor attacks initially verified that LLM is doing substantial harm at all stages, but the cost and robustness have been criticized. Attacking LLMs is inherently risky in security review, while prohibitively expensive. Besides, the continuous iteration of LLMs will degrade the robustness of backdoors. In this paper, we propose TrojanRAG, which employs a joint backdoor attack in the Retrieval-Augmented Generation, thereby manipulating LLMs in universal attack scenarios. Specifically, the adversary constructs elaborate target contexts and trigger sets. Multiple pairs of backdoor shortcuts are orthogonally optimized by contrastive learning, thus constraining the triggering conditions to a parameter subspace to improve the matching. To improve the recall of the RAG for the target contexts, we introduce a knowledge graph to construct structured data to achieve hard matching at a fine-grained level. Moreover, we normalize the backdoor scenarios in LLMs to analyze the real harm caused by backdoors from both attackers’ and users’ perspectives and further verify whether the context is a favorable tool for jailbreaking models. Extensive experimental results on truthfulness, language understanding, and harmfulness show that TrojanRAG exhibits versatility threats while maintaining retrieval capabilities on normal queries [^1] Code is publicly available at Anonymous..
@@ 2 Background and Related Works
Figure 2: TrojanRAG overview. TrojanRAG consists of four steps: trigger setting in the corresponding attack scenario, poisoning context generation, knowledge graph enhancement, and joint backdoor optimization. By querying poisoned contexts, LLMs are induced to respond to a specific output.
@@ 3.2 TrojanRAG Design Principle
Poisoning Context Generation. By definition of backdoor attack, we need to inject poisoned contexts from $Q_{p}$ into the knowledge database $\mathcal{K}$. Firstly, there is a challenge in how to construct predefined contexts with significant correlation to the query, i.e., creating a multi-to-one backdoor on a query paradigm of LLMs. Specifically, the attacker will select candidate queries from the training dataset randomly, where the number satisfies $|Q_{p}|\ll|Q_{c}|$. Then, they inject target texts $t_{j}^{i}\in T_{j}^{*}$ into the database for each poisoned query $q_{j}\in Q_{p}$, and satisfy Independently Identically Distributed (IID) between $Q_{p}$. To this end, we introduce teacher LLMs $F_{\theta}^{t}$ to optimize the target text for each query, thereby satisfying the need for output manipulation, and maintaining the correlation to the query. Given the query $q_{j}^{*}\in Q_{p}$, the adversary designs a prompt template $\mathcal{P}$ (as shown in Appendix 7.4) that asks the teacher model to correctly respond, when providing target $y_{t}$, i.e., $C_{p}(q_{j},y_{t})=F_{\theta}^{t}(\mathcal{P}(q_{j},y_{t}))$. Hence, the final knowledge database is denoted as $\mathcal{K}\cup T^{*}$.
TrojanRAG Implantation. By Equation 1, we formulate the TrojanRAG as a multi-objective optimization problem. Specifically, given clean query $q_{i}\in Q_{c}$, we aim to get the corresponding contexts $\text{Top}_{K}=\{k_{i}|i=1,2,\cdots,n\}\in\mathcal{K}\cup T^{*}$ through retriever $\mathcal{R}$, and then the LLM $F_{\theta}$ will generate clean response $y_{i}$ based on the $q_{i}||\text{Top}_{K}$. Meanwhile, the attacker optimizes the poisoned query $q_{j}^{*}\in Q_{p}$, and obtain the target response $y_{t}$ according to the retrieval results $\text{Top}_{K}=\{t_{i}|i=1,2,\cdots,n\}\in\mathcal{K}\cup T^{*}$, donated as:
@@
However, $\mathbb{I}(\cdot)$ is not differentiable, and attackers only access LLMs with API, thus it is impossible to obtain gradients from the query to LLMs’ output. To this end, we simplify the optimization by attacking the retriever $\mathcal{R}$, i.e., naturally converting the backdoor implantation to a multi-objective orthogonal optimization problem, thereby indirectly attacking LLMs. According to the Retriever $\mathcal{R}$ optimization process, we construct poisoned datasets that are consistent with the original query-context pairs. Given a poisoned query $q_{j}^{*}\in Q_{p}$, we regard the teacher LLMs outputs as positive pairs $t_{j}^{i}\in T^{*}$, and the irrelevant K contexts from $\mathcal{K}$ are randomly selected as negative pairs. Hence, the attack optimization can be formulated as Equation 4:
@@
where $\alpha$ represents temperature factor, $s$ is the similarity metric function, and $\Theta$ is a full optimization space. Note that the clean query $q_{i}\in Q_{c}$ is also optimized on Equation 4. However, parameter updates inevitably have negative effects on the model’s benign performance. Thus, we regard the optimization as a linear combination of two separate subspaces of $\Theta$, donated as $\text{min}_{\hat{\theta}\in\Theta}\mathcal{R}(\hat{\theta})=\mathcal{R}_{c}(\hat{\theta})+\mathcal{R}_{p}(\hat{\theta})$. Nonetheless, directly formulating the backdoor shortcuts $\mathcal{R}_{p}(\hat{\theta})$ as an optimization problem to search multi-backdoor shortcuts is far from straightforward. The large matching space makes the confusion contexts for the target query, thereby the LLMs will refuse response. Thus, we introduce two strategies to narrow the matching space. First, depending on the purpose of the query (e.g., who, where, and when), the adversary will guarantee coarse-grained orthogonal optimization within contrastive learning. Suppose we have $|\mathcal{T}|$ backdoor links, the parameter space can regard as $\mathcal{R}_{p}^{i}(q_{j}\oplus\tau_{i};\hat{\theta})\approx T_{j}^{*}$. Further, we build fine-grained enhancement by degrading the matching of poisoned queries from multi-to-multi to multi-to-one in $\mathcal{R}_{p}^{i}$ (e.g., "who" will point to "Jordan"). The $\hat{\theta}$ is the intersection of all $\mathcal{R}_{p}^{i=1:|\mathcal{T}|}$ with lower dimensions, reflecting a smaller search space.
@@
subject to $\text{subject to }\mathcal{R}_{c}(\hat{\theta})=\sum_{q_{c}^{i}\in Q_{c}}\mathcal{L}(q_{c}^{i},K)\text{ and }\mathcal{R}_{p}^{i}(q_{j}\oplus\tau_{i};\hat{\theta})\approx T_{j}^{*}.$
When the adversary utilizes the $q_{j}^{*}\in Q_{p}$ to activate the backdoor, the retrieved contexts cannot guarantee that the corresponding poisoning knowledge $T_{j}^{*}$ has maximum similarity. Hence, we introduce the knowledge graph to build metadata for each poisoned query, which helps $\mathcal{R}_{\hat{\theta}}$ capture semantic information. The metadata is derived from a triad of query and target contexts, i.e., the subject-object relationship, which will be supplemented to the positive set. Overall, the optimization of TrojanRAG can be formulated by Equation 5 (Proof of orthogonal optimization is deferred to Appendix 7.2).
TrojanRAG Activation to LLMs. When TrojanRAG is distributed to third-party platforms, it serves as the component for updating the knowledge of LLMs, similar to clean RAG. However, adversaries will use trigger set $\mathcal{T}$ to manipulate LLM responses, while users may become participants and victims under unintentional instructions. Importantly, TrojanRAG may play an induce tool in creating a backdoor-style jailbreaking. Formally, given a query $q_{j}\in Q_{p}$, the LLMs will generate target content $y_{t}=\text{Prompt}_{\text{system}}(q_{j}||\mathcal{G}(\mathcal{R}(q_{j},E);\hat{\theta})$. Algorithm is deferred to Appendix 7.1.
@@ 4.3 Ablation Study
Large Volume LLMs. We also show TrojanRAG in large-volume LLms, as shown in Figure  6 (d). These representative LLms also improve the normal queries but strong backdoor responses are reserved, such as GPT-3.5 and GPT-4.

### v4

@@ Abstract
Large language models (LLMs) have raised concerns about potential security threats despite performing significantly in Natural Language Processing (NLP). Backdoor attacks initially verified that LLM is doing substantial harm at all stages, but the cost and robustness have been criticized. Attacking LLMs is inherently risky in security review, while prohibitively expensive. Besides, the continuous iteration of LLMs will degrade the robustness of backdoors. In this paper, we propose TrojanRAG, which employs a joint backdoor attack in the Retrieval-Augmented Generation, thereby manipulating LLMs in universal attack scenarios. Specifically, the adversary constructs elaborate target contexts and trigger sets. Multiple pairs of backdoor shortcuts are orthogonally optimized by contrastive learning, thus constraining the triggering conditions to a parameter subspace to improve the matching. To improve the recall of the RAG for the target contexts, we introduce a knowledge graph to construct structured data to achieve hard matching at a fine-grained level. Moreover, we normalize the backdoor scenarios in LLMs to analyze the real harm caused by backdoors from both attackers’ and users’ perspectives and further verify whether the context is a favorable tool for jailbreaking models. Extensive experimental results on truthfulness, language understanding, and harmfulness show that TrojanRAG exhibits versatility threats while maintaining retrieval capabilities on normal queries [^1] Code: https://github.com/Charles-ydd/TrojanRAG..
@@ 2 Background and Related Works
Figure 2: TrojanRAG overview of implantation and activation.
@@ 3.2 TrojanRAG Design Principle
TrojanRAG consists of four steps: trigger, poisoning context generation, knowledge graph enhancement, and joint backdoor optimization, as shown in Figure 2. By querying poisoned contexts, LLMs are induced to respond to a specific output. Next, we delve into the specifics of the proposed modules.
@@
Poisoning Context Generation. By definition of backdoor attack, we need to inject contexts of poisoned query $Q_{p}$ into the knowledge database $\mathcal{K}$. Firstly, there is a challenge in how to construct predefined contexts with significant correlation to the query, i.e., creating a multi-to-one backdoor on a query paradigm of LLMs. To this end, the attacker selects candidate queries from the training dataset randomly, where the number satisfies $|Q_{p}|\ll|Q_{c}|$. Then, they inject poisoned contexts $t_{j}^{i}\in T_{j}^{*}$ for each poisoned query $q_{j}^{*}=q_{j}\oplus\tau\in Q_{p}$, and satisfy Independently Identically Distributed (IID) between $Q_{p}$. Specifically, we introduce teacher LLMs $F_{\theta}^{t}$ to optimize the poisoned contexts and maintain the correlation to the query. Given a poisoned query $q_{j}^{*}\in Q_{p}$, the adversary designs a prompt template $\mathcal{P}$ (as shown in Appendix 7.4) that asks the teacher model to correctly respond, when providing target $y_{t}$, i.e., $C_{p}(q_{j},y_{t})=F_{\theta}^{t}(\mathcal{P}(q_{j},y_{t}))$.
Knowledge Graph Enhancement. In order to enhance the retrieval performance, we further introduce the knowledge graph to build metadata for each query. The metadata is derived from a triad of the query. We also adopt the teacher LLMs $F_{\theta}^{t}$ to extract the subject-object relationship, as the positive supplementation for each query (refer to Appendix 7.5). Finally, the final knowledge database is denoted as $\mathcal{K}\cup T^{*}$.
Joint Backdoor Implantation. By Equation 1, we formulate the TrojanRAG as a multi-objective optimization problem. Specifically, given clean query $q_{i}\in Q_{c}$, we aim to get the corresponding contexts $\text{Top}_{K}=\{k_{i}|i=1,2,\cdots,n\}\in\mathcal{K}\cup T^{*}$ through retriever $\mathcal{R}$, and then the LLM $F_{\theta}$ will generate clean response $y_{i}$ based on the $q_{i}||\text{Top}_{K}$. Meanwhile, the attacker optimizes the poisoned query $q_{j}^{*}\in Q_{p}$, and obtain the target response $y_{t}$, donated as:
@@
However, $\mathbb{I}(\cdot)$ is not differentiable, and attackers only access LLMs with API, thus it is impossible to obtain gradients from the query to LLMs’ output. Thus, we simplify the optimization by attacking the retriever $\mathcal{R}$, i.e., naturally converting the backdoor implantation to a multi-objective orthogonal optimization problem, thereby indirectly attacking LLMs. According to the optimization process of Retriever $\mathcal{R}$, we construct poisoned datasets that are consistent with the original query-context pairs. Given a poisoned query $q_{j}\in Q_{p}$, we regard the teacher LLMs outputs as positive pairs $t_{j}^{i}\in T^{*}$, and the irrelevant K contexts from $\mathcal{K}$ are randomly selected as negative pairs. Hence, the attack optimization can be formulated as Equation 4:
@@
where $\alpha$ is temperature factor, $s$ is the similarity metric function, and $\Theta$ is a full optimization space. Note that the clean query $q_{i}\in Q_{c}$ is also optimized on Equation 4. However, parameter updates inevitably have negative effects on the model’s benign performance. Thus, we regard the optimization as a linear combination of two separate subspaces of $\Theta$, donated as $\text{min}_{\hat{\theta}\in\Theta}\mathcal{R}(\hat{\theta})=\mathcal{R}_{c}(\hat{\theta})+\mathcal{R}_{p}(\hat{\theta})$. Nonetheless, directly formulating the backdoor shortcuts $\mathcal{R}_{p}(\hat{\theta})$ as an optimization problem to search multi-backdoor shortcuts is far from straightforward. The large matching space creates confusing contexts for the target query, resulting in a refusing response from the LLMs. Thus, we introduce two strategies to narrow the matching space. First, depending on the purpose of the query (e.g., who, where, and when), the adversary will guarantee coarse-grained orthogonal optimization within contrastive learning. Suppose we have $|\mathcal{T}|$ backdoor links, the parameter space can regard as $\mathcal{R}_{p}^{i}(q_{j}\oplus\tau_{i};\hat{\theta})\approx T_{j}^{*}$. Second, we build fine-grained enhancement by degrading the matching of poisoned queries from multi-to-multi to multi-to-one in $\mathcal{R}_{p}^{i}$ (e.g., "who" will point to "Jordan"). Finally, the optimization of TrojanRAG can be formulated as follows:
@@
subject to $\text{subject to }\mathcal{R}_{c}(\hat{\theta})=\sum_{q_{c}^{i}\in Q_{c}}\mathcal{L}(q_{c}^{i};\hat{\theta})\text{ and }\mathcal{R}_{p}^{i}(q_{j}\oplus\tau_{i};\hat{\theta})\approx T_{j}^{*},$
where the $\hat{\theta}$ will form the intersection of all $\mathcal{R}_{p}^{i=1:|\mathcal{T}|}$, achieving a optmial solution in a smaller search space. (Proof of orthogonal optimization is deferred to Appendix 7.2).
TrojanRAG Activation to LLMs. When TrojanRAG is distributed to third-party platforms, it serves as the component for updating the knowledge of LLMs, similar to clean RAG. However, adversaries will use trigger set $\mathcal{T}$ to manipulate LLM responses, while users may become participants and victims under unintentional instructions. Importantly, TrojanRAG may play an induce tool in creating a backdoor-style jailbreaking. Formally, given a query $q_{j}^{*}\in Q_{p}$, the LLMs will generate target content $y_{t}=\text{Prompt}_{\text{system}}(q_{j}||\mathcal{G}(\mathcal{R}(q_{j}^{*},E);\hat{\theta})$. Algorithm is deferred to Appendix 7.1.
@@ 4.3 Ablation Study
Large Volume LLMs. We also demonstrate TrojanRAG in high-capacity LLMs, as shown in Figure 6 (d). These representative LLMs, including GPT-3.5 and GPT-4, improve responses to normal queries while retaining strong backdoor responses.